A certain amount of a metal whose equivalent mass is 28 displaces 0.7L of hydrogen gas at S.T.P from an acid hence mass of the element is

Grade:12th pass

2 Answers

Vikas TU
7 years ago
moles at STP
n = V/22.4 -=> 0.4=7/22.4 = > 0.3125 moles.
Now, Weight of element = 0.3125 * 28 * 2 
=> 1.7583 gms approx.
Bhavya
7 years ago
Dear student
Here we have to apply law of equivalence,
equivalent of metal = equivalent of hydrogen
mass/Equivalent mass of metal = Volume of H2/11.2
mass/28 = 0.7/11.2
mass = 1.75 g